Find T,N,B if possible for the following curves
(a)r(t)=(t³/3)i+(t²/2)j,t>0
(b) r(t)=(t³/3)I+(t²/2)j,t>=0
1
Expert's answer
2020-12-10T14:08:17-0500

The unit tangent vector "\overrightarrow{T}" , the unit normal vector "\overrightarrow{N}"  and the unit binormal vector "\overrightarrow{B}"  are three mutually perpendicular vectors used to describe a curve in two or three dimensions. This moving coordinate system is attached to the curve and describes the shape of the curve independent of any parameterization.

If the curve is given parametrically by



"x=x(t)\quad\quad y=y(t)"

the position, unit tangent, unit normal, and unit binormal vectors are



"\overrightarrow{r}(t)=x(t)\cdot \overrightarrow{i}+y(t)\cdot \overrightarrow{j}\\[0.3cm]
\overrightarrow{T}=\frac{\overrightarrow{r}'(t)}{\left|\overrightarrow{r}'(t)\right|} \\[0.3cm]
\overrightarrow{N}=\frac{\overrightarrow{T}'(t)}{\left|\overrightarrow{T}'(t)\right|} \\[0.3cm]
\overrightarrow{B}=\overrightarrow{T}(t)\times\overrightarrow{N}\\[0.3cm]"

1 case :



"\overrightarrow{r}(t)=\frac{t^3}{3}\cdot\overrightarrow{i}+\frac{t^2}{2}\cdot\overrightarrow{j},\,\,\,t>0\\[0.3cm]
\overrightarrow{r}'(t)=\frac{d}{dt}\left(\frac{t^3}{3}\cdot\overrightarrow{i}+\frac{t^2}{2}\cdot\overrightarrow{j}\right)=\frac{3t^2}{3}\cdot\overrightarrow{i}+\frac{2t}{2}\cdot\overrightarrow{j}\\[0.3cm]
\boxed{\overrightarrow{r}'(t)=t^2\cdot\overrightarrow{i}+t\cdot\overrightarrow{j}}\\[0.3cm]
\left|\overrightarrow{r}'(t)\right|=\sqrt{(t^2)^2+t^2}=\sqrt{t^2\cdot\left(t^2+1\right)}=|t|\cdot\sqrt{t^2+1}\\[0.3cm]
\left|\overrightarrow{r}'(t)\right|=[t>0]=t\cdot\sqrt{t^2+1}\\[0.3cm]
\overrightarrow{T}(t)=\frac{\overrightarrow{r}'(t)}{\left|\overrightarrow{r}'(t)\right|}=\frac{t^2\cdot\overrightarrow{i}+t\cdot\overrightarrow{j}}{t\cdot\sqrt{t^2+1}}=\frac{t^2\cdot\overrightarrow{i}}{t\cdot\sqrt{t^2+1}}+\frac{t\cdot\overrightarrow{j}}{t\cdot\sqrt{t^2+1}}\\[0.3cm]
\boxed{\overrightarrow{T}(t)=\frac{t\cdot\overrightarrow{i}}{\sqrt{t^2+1}}+\frac{\overrightarrow{j}}{\sqrt{t^2+1}}}\\[0.3cm]
\overrightarrow{T}'(t)=\frac{d}{dt}\left(\frac{t\cdot\overrightarrow{i}}{\sqrt{t^2+1}}+\frac{\overrightarrow{j}}{\sqrt{t^2+1}}\right)=\\[0.3cm]
=\left(\frac{1}{\sqrt{t^2+1}}+t\cdot\left(-\frac{1}{2}\cdot\frac{2t}{\left(t^2+1\right)^{3/2}}\right)\right)\cdot\overrightarrow{i}+\left(-\frac{1}{2}\cdot\frac{2t}{\left(t^2+1\right)^{3/2}}\right)\cdot\overrightarrow{j}=\\[0.3cm]
=\left(\frac{t^2+1-t^2}{\left(t^2+1\right)^{3/2}}\right)\cdot\overrightarrow{i}+\left(-\frac{t}{\left(t^2+1\right)^{3/2}}\right)\cdot\overrightarrow{j}=\\[0.3cm]
=\left(\frac{1}{\left(t^2+1\right)^{3/2}}\right)\cdot\overrightarrow{i}+\left(-\frac{t}{\left(t^2+1\right)^{3/2}}\right)\cdot\overrightarrow{j}\\[0.3cm]
\left|\overrightarrow{T}(t)\right|=\sqrt{\left(\frac{1}{\left(t^2+1\right)^{3/2}}\right)^2+\left(\frac{t}{\left(t^2+1\right)^{3/2}}\right)^2}=\\[0.3cm]
=\sqrt{\frac{t^2+1}{\left(t^2+1\right)^3}}=\sqrt{\frac{1}{\left(t^2+1\right)^2}}=\frac{1}{t^2+1}\longrightarrow\frac{1}{\left|\overrightarrow{T}'(t)\right|}=(t^2+1)\\[0.3cm]
\overrightarrow{N}(t)=\frac{\overrightarrow{T}'(t)}{\left|\overrightarrow{T}'(t)\right|}=(t^2+1)\cdot\left(\left(\frac{1}{\left(t^2+1\right)^{3/2}}\right)\cdot\overrightarrow{i}+\left(-\frac{t}{\left(t^2+1\right)^{3/2}}\right)\cdot\overrightarrow{j}\right)\\[0.3cm]
=\left(\frac{1}{\left(t^2+1\right)^{1/2}}\right)\cdot\overrightarrow{i}+\left(-\frac{t}{\left(t^2+1\right)^{1/2}}\right)\cdot\overrightarrow{j}\\[0.3cm]
\boxed{\overrightarrow{N}(t)=\frac{\overrightarrow{i}}{\sqrt{t^2+1}}-\frac{t\cdot\overrightarrow{j}}{\sqrt{t^2+1}}}\\[0.3cm]
\overrightarrow{B}=\overrightarrow{T}\times\overrightarrow{N}=\left|\begin{array}{ccc}
\overrightarrow{i}&\overrightarrow{j}&\overrightarrow{k}\\
T_x&T_y&T_z=0\\
N_x&N_y&N_z=0
\end{array}\right|=\overrightarrow{k}\cdot(T_xN_y-T_yN_x)=\\[0.3cm]
=\overrightarrow{k}\cdot\left(\left(\frac{t}{\sqrt{t^2+1}}\right)\cdot\left(\frac{-t}{\sqrt{t^2+1}}\right)-\left(\frac{1}{\sqrt{t^2+1}}\right)\cdot\left(\frac{1}{\sqrt{t^2+1}}\right)\right)=\\[0.3cm]
=\overrightarrow{k}\cdot\left(\frac{-(t^2+1)}{t^2+1}\right)=-1\cdot\overrightarrow{k}\\[0.3cm]
\boxed{\overrightarrow{B}=-1\cdot\overrightarrow{k}}"

2 case :



"\overrightarrow{r}(t)=\frac{t^3}{3}\cdot\overrightarrow{i}+\frac{t^2}{2}\cdot\overrightarrow{j},\,\,\,t\ge0\\[0.3cm]
\overrightarrow{r}'(t)=\frac{d}{dt}\left(\frac{t^3}{3}\cdot\overrightarrow{i}+\frac{t^2}{2}\cdot\overrightarrow{j}\right)=\frac{3t^2}{3}\cdot\overrightarrow{i}+\frac{2t}{2}\cdot\overrightarrow{j}\\[0.3cm]
\boxed{\overrightarrow{r}'(t)=t^2\cdot\overrightarrow{i}+t\cdot\overrightarrow{j}}\\[0.3cm]
\left|\overrightarrow{r}'(t)\right|=\sqrt{(t^2)^2+t^2}=\sqrt{t^2\cdot\left(t^2+1\right)}=|t|\cdot\sqrt{t^2+1}\\[0.3cm]
\left|\overrightarrow{r}'(t)\right|=[t\ge0]=t\cdot\sqrt{t^2+1}\\[0.3cm]"

There is a problem with the parameter value "t=0" , since



"\overrightarrow{r}'(0)=0^2\cdot\overrightarrow{i}+0\cdot\overrightarrow{j}\\[0.3cm]
\left|\overrightarrow{r}'(0)\right|=0\cdot\sqrt{0^2+1}=0\\[0.3cm]
\overrightarrow{T}(0)=\frac{0}{0}\cdot\overrightarrow{i}+\frac{0}{0}\cdot\overrightarrow{j}-\text{formal entry}"

It means that "\nexists\overrightarrow{T}(0)" and as a consequence "\nexists\overrightarrow{N}(0)" and "\nexists\overrightarrow{B}(0)" .


ANSWER



"\overrightarrow{r}(t)=\frac{t^3}{3}\cdot\overrightarrow{i}+\frac{t^2}{2}\cdot\overrightarrow{j},\,\,\,t>0\\[0.3cm]
\overrightarrow{T}(t)=\frac{t\cdot\overrightarrow{i}}{\sqrt{t^2+1}}+\frac{\overrightarrow{j}}{\sqrt{t^2+1}}\\[0.3cm]
\overrightarrow{N}(t)=\frac{\overrightarrow{i}}{\sqrt{t^2+1}}-\frac{t\cdot\overrightarrow{j}}{\sqrt{t^2+1}}\\[0.3cm]
\overrightarrow{B}=-1\cdot\overrightarrow{k}\\[0.3cm]
\overrightarrow{r}(t)=\frac{t^3}{3}\cdot\overrightarrow{i}+\frac{t^2}{2}\cdot\overrightarrow{j},\,\,\,t\ge0\\[0.3cm]
\overrightarrow{T}(t)=\frac{t\cdot\overrightarrow{i}}{\sqrt{t^2+1}}+\frac{\overrightarrow{j}}{\sqrt{t^2+1}},\,\,\,t>0\,\,\,\text{and}\,\,\,\nexists\overrightarrow{T}(0)\\[0.3cm]
\overrightarrow{N}(t)=\frac{\overrightarrow{i}}{\sqrt{t^2+1}}-\frac{t\cdot\overrightarrow{j}}{\sqrt{t^2+1}},\,\,\,t>0\,\,\,\text{and}\,\,\,\nexists\overrightarrow{N}(0)\\[0.3cm]
\overrightarrow{B}=-1\cdot\overrightarrow{k},\,\,\,t>0\,\,\,\text{and}\,\,\,\nexists\overrightarrow{B}(0)"